# Macroeconomics of Racial Disparities: Discrimination, Labor Market, and Wealth

### v2

@@ 6 Conclusion
This study examines the effects of racial hiring discrimination within a frictional labor market on employment, wage disparities, and wealth accumulation between black and white workers in the U.S. We develop a search-and-matching model incorporating firms with and without racial prejudices alongside race-specific pathways for wealth accumulation. Our findings reveal that racial discrimination in hiring significantly exacerbates wage, unemployment, and wealth gaps in the steady state and business cycles. Moreover, discriminatory hiring disproportionately consigns black workers to the lower end of the wealth spectrum. Contrary to conventional discrimination theories, our analysis suggests that discriminatory hiring persists as an equilibrium outcome within frictional markets. Eliminating these discriminatory factors results in a decrease in overall welfare. Our findings shed light on the enduring nature of Black-White racial disparities in the U.S., offering insights into the interplay between discrimination, labor market dynamics, and wealth accumulation.
@@ Appendix B Comparing the impact of bargaining power and extreme wealth shock
Table B.1 compares the benchmark model to models that remove non-market disparities, particularly in equalizing racial bargaining power and extreme wealth shocks. Equalizing black workers’ bargaining power with that of white workers (Column 2, $\xi_{bl}=\xi_{wh}$) depresses black workers’ job-finding rate but directly raises their bargained wage outcomes. This is because non-prejudiced firms retain less profit, hence post fewer vacancies. On the net, it also reduces white workers’ job-finding rate and average wage rate. As a result, the average wage gap shrinks to 19% (with a wage ratio of 81%), explaining about 24% of the racial wage gap.
@@
job finding rate - black 0.50 0.43 0.48
@@
mean wage - black 1.60 1.71 1.66
mean wage - white 2.13 2.12 2.15
mean wage ratio 0.75 0.81 0.77
mean wealth ratio 0.28 0.29 0.77
@@
$p$ firm profit 0.04 0.04 0.04
np firm profit 0.03 0.02 0.02
@@
np firm vacancy 0.05 0.04 0.05
p firm employment 0.54 0.58 0.56
np firm employment 0.40 0.35 0.38
@@
$p$ market tightness 0.40 0.46 0.43
$np$ market tightness 0.73 0.60 0.69
@@
Y 3.06 3.05 3.08
K/Y 10.26 10.25 10.45
average wage 2.04 2.05 2.06
unemp rate 0.07 0.07 0.07
@@
In Column 3 ($\epsilon_{bl}=\epsilon_{wh}$), we assign black workers the same conditions in accumulating wealth as white workers. The effect resembles assigning higher bargaining power to black workers. This is due to the importance of personal wealth in partial self-insurance against uncertain adverse outcomes (Nakajima, 2012). With a lower probability of losing their wealth, black workers can accumulate more personal wealth. Higher personal wealth gives black workers a higher reservation value when bargaining with firms. Effectively, black workers can negotiate for better wage outcomes (1.66 compared to 1.60 in the benchmark model). Similar to directly assigning higher bargaining power, the resulting job-finding rates decrease, though to a smaller extent. The smaller extreme wealth destruction rate also translates into better aggregate capital accumulation and increased aggregate output. It spills over to an increase in the average wage rate for all. Though it has a much lesser impact on the unemployment rate gap, it raises the average black-to-white wage ratio to 77% and reduces the wage gap to 23%, explaining 8% of the racial wage gap. Since it directly raises the wealth position of black workers, the mean and median black-to-white wealth ratios increase drastically (77% and 88% from the benchmark 28% and 33%).
Appendix C Impulse responses
@@
We follow Boppart et al. (2018) to simulate the business cycle moments of the model from the derivatives of the impulse response functions. For the impulse response functions to represent the numerical derivative, the magnitude of the MIT shock should be small. This method also requires the business cycle dynamics to be well approximated as a linear system. Figure C.1 presents the impulse response functions for the benchmark model to a 1% positive and negative TFP shock at date 0. The shock gradually returns to a steady state with a persistence of 0.95. Our model produces symmetric impulse responses across all variables.
Figure C.1: Comparing impulse responses

### v3

@@ 6 Conclusion
This study examines the effects of discrimination in the hiring of black workers on employment, wage disparities, and wealth accumulation in the US. We develop a search-and-matching model featuring firms with and without racial prejudices. Our findings reveal that racial discrimination in hiring significantly exacerbates wage, unemployment, and wealth gaps in the steady state and business cycles. In addition, discriminatory hiring places black workers disproportionately at the lower end of the wealth spectrum. Contrary to conventional discrimination theories, our analysis suggests that discriminatory hiring persists as an equilibrium outcome in frictional labor markets. Further, eliminating discriminatory hiring results in an increase in overall economic welfare. Our findings shed light on the enduring nature of black-white racial disparities in the US, offering insights into the interplay between discrimination, labor market dynamics, and wealth accumulation.
@@ Appendix B Comparing the impact of bargaining power and extreme wealth shock
Our model separately calibrates racial differences in bargaining power and the likelihood of extreme wealth destruction - features we refer to as non-market racial disparities. We conduct additional counterfactual analyses to assess the impact of these non-market factors relative to hiring discrimination. Table B.1 compares the benchmark model to versions that eliminate non-market disparities by equalizing bargaining power and extreme wealth shocks across racial groups. When we set the bargaining power of black workers equal to that of white workers (Column 2, $\xi_{bl}=\xi_{wh}$), their bargained wages rise directly. However, the job-finding rate for black workers falls. This outcome occurs because non-prejudiced firms retain lower profits and, therefore, post fewer vacancies, which also reduces the job-finding rate and average wage for white workers. Consequently, the average wage gap shrinks to 20% (a wage ratio of 80%), accounting for roughly 20% of the observed racial wage gap.
@@
job finding rate - black 0.49 0.44 0.48
@@
mean wage - black 1.59 1.70 1.65
mean wage - white 2.12 2.11 2.13
mean wage ratio 0.75 0.80 0.77
mean wealth ratio 0.29 0.30 0.77
@@
$p$ firm profit 0.04 0.05 0.04
np firm profit 0.02 0.02 0.02
@@
np firm vacancy 0.05 0.04 0.04
p firm employment 0.57 0.61 0.58
np firm employment 0.37 0.33 0.35
@@
$p$ market tightness 0.39 0.44 0.42
$np$ market tightness 0.71 0.60 0.67
@@
Y 3.07 3.06 3.08
K/Y 10.26 10.25 10.40
average wage 2.05 2.05 2.06
unemp rate 0.06 0.06 0.06
@@
In Column 3 ($\epsilon_{bl}=\epsilon_{wh}$), we assign black workers the same extreme wealth shock as white workers, an external condition governing the accumulating wealth. The effect resembles assigning higher bargaining power to black workers. This is due to the importance of personal wealth in helping workers self-insurance against uncertain adverse outcomes (Nakajima, 2012). With a lower probability of losing their wealth, black workers can accumulate more personal wealth. Higher personal wealth gives black workers a higher reservation value when bargaining with firms. Effectively, black workers can negotiate for better wage outcomes (1.65 compared to 1.59 in the benchmark model). Similar to having higher bargaining power, the resulting job-finding rates decrease, though to a smaller extent. The lower extreme wealth destruction rate also translates to higher aggregate capital accumulation and increased aggregate output. It spills over to an increase in the average wage rate for all, raising the average black-white wage ratio to 77%, thus explaining around 8% of the racial wage gap. Since it directly raises the wealth position of black workers, the mean and median black-white wealth ratios increase drastically (77% and 88% from the benchmark levels of 29% and 33%).
Appendix C Welfare calculation
We follow Krusell et al. (2010) in measuring the welfare effects of discriminatory hiring on black and white workers. We calculate the consumption equivalence in a counterfactual scenario without $p$ firms. Under the benchmark model, let $V(e,R,s,a)=\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log(c_{t})$ be the maximal value of the individual with employment status $e$, race $R$, productivity $s$, and asset $a$. In an alternative economy, let $\tilde{V}(e,R,s,a)=\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log(\tilde{c}_{t})$ be the maximal value of individuals with each corresponding state. We examine the welfare change between the two economies through consumption equivalence $\Omega$, following the equation: $\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log((1+\Omega)c_{t})=\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log(\tilde{c}_{t})$.
Under $\log$ utility, we derive $\Omega=\exp((\tilde{V}-V)(1-\beta))-1$. We aggregate the individual-level consumption equivalence, $\Omega$s, using the distribution of the counterfactual economy to calculate the average welfare change. We aggregate over the counterfactual distribution rather than the benchmark distribution because eliminating $p$ firms in the model removes the distribution of white workers from the states associated with $p$ firms. Aggregating using the benchmark distribution overstates the welfare change for white workers without accounting for this distribution shift.
Appendix D Measuring the business cycle properties for wage
Aggregate data can be misleading when estimating wage cyclicality because shifts in worker composition over the business cycle tend to overweight lower-skilled (and lower-wage) workers during expansions, resulting in countercyclical bias. To overcome this issue, we use individual-level PSID data (1996-2014), which allows us to hold the composition of workers fixed over time.
Following the methodology of Bils (1985), Solon et al. (1994), and Devereux (2001), we estimate wage cyclicality by regressing the change in individual log wages on the change in log GDP while controlling for key demographic characteristics. Our specification is as follows:
$\Delta\log w_{it}=\alpha+\beta\Delta\log Y_{t}+\gamma x_{it}+\nu_{t}+\upsilon_{s}+\epsilon_{it},$ (D.1)
where $Y_{t}$ is real GDP. The demographic variables $x_{it}$ include gender, education, labor market experience, and marital status. We also include year fixed effects, $\nu_{t}$, and state fixed effects, $\upsilon_{s}$.
In this regression, the coefficient $\beta$ captures the cyclicality of wages. We then compute wage volatility as the standard deviation of wage growth after controlling for the demographics. The year fixed effects from the regression reflect aggregate wage growth and provide a measure of this volatility. This approach ensures that our estimates of wage cyclicality and volatility reflect true economic fluctuations rather than changes in workforce composition.
Appendix E Impulse responses
@@
We follow Boppart et al. (2018) to simulate the business cycle moments of the model from the derivatives of the impulse response functions. For the impulse response functions to represent the numerical derivative, the magnitude of the MIT shock should be small. This method also requires the business cycle dynamics to be well approximated as a linear system. Figure E.1 presents the impulse response functions for the benchmark model to a 1% positive and a 1% negative TFP shock at date 0. The shock gradually returns to a steady state with a persistence of 0.95. Our model produces symmetric impulse responses for positive and negative TFP shocks for all variables.
Figure E.1: Comparing impulse responses